OpenStudy (cutestgirlaround):

6h−(9−3h)=6−3(h+2) A. h=3/2 <--- B. h=34 C. h=7/2 <--- D. h=17/6

OpenStudy (cutestgirlaround):

@Qwertty123

OpenStudy (qwertty123):

Okay soo... Simplify both sides of the equation. 6h−(9−3h)=6−3(h+2)

OpenStudy (qwertty123):

Then.. Add 3h to both sides. 9h−9+3h=−3h+3h 12h−9=0

OpenStudy (qwertty123):

Then...Add 9 to both sides. 12h−9+9=0+9 12h=9

OpenStudy (qwertty123):

Then.. Divide both sides by 12. To get you answer...

OpenStudy (cutestgirlaround):

oh so Its B 3/4

OpenStudy (qwertty123):

Yes!

OpenStudy (cutestgirlaround):

thanks!!! I have 3 more questions would you like to help me out on those ?

OpenStudy (qwertty123):

Sure open up a new one..

OpenStudy (cutestgirlaround):

ok thanks

OpenStudy (qwertty123):

No problem